# Constraints on the progenitor and explosion of SN 2024ggi in harmony with pre-explosion detection and hydrodynamic simulations

### v1

@@ 2.3 Constrain on the physical properties of the possible progenitor
The evolutionary tracks of the models shown in Figure 2 start from ZAMS and terminate at their respective pre-SN stages marked by for each model. The location of the progenitor detected in the pre-explosion images from HST (results from D. Xiang et al., 2024) is also indicated. The effective temperature , and luminosity log(/L⊙) of the 11 M⊙ model at pre-SN stage align very well with the progenitor directly detected from HST in archival images. At the pre-SN stage, the 11 M⊙ has a radius of 800 R⊙. Additionally, the age (log[/yr]) of the 11 M⊙ model at the pre-SN stage is 7.38, matching closely with X. Hong et al. (2024).
Although the pre-SN properties of 11 M⊙ model match well with the progenitor detected in pre-explosion images, it is difficult to completely rule out the 10 M⊙ and 12 M⊙ models. These models have pre-SN radii of 690 R⊙ and 900 R⊙, respectively. Thus, for the progenitor of SN 2024ggi, we provide a constrain of 11 M⊙ on the initial mass, and a range of pre-SN radius of 690 R⊙–900 R⊙. These constraints align nicely with the values obtained by analyzing the pre-explosion images (as in D. Xiang et al., 2024; X. Hong et al., 2024). The error bar on our initial mass constraint represents the range of masses lying in the closest vicinity of the error bar of the pre-explosion detected progenitor on the HR-diagram. The details of the physical properties of the models at different stages are provided in Table 1.
@@ 3 Constraints on the explosion properties
In the previous section, we explored the physical properties of the models at their pre-SN stages marked by the onset of Fe-core infall. Next, the models are allowed to continue evolving until they reach near-shock breakout (SBO) stages utilizing the ccsn_IIp test-suite directory of MESA. The pre-SN model from the 12M_pre_ms_to_core_collapse directory is incorporated as the starting point in the ccsn_IIp test-suite directory. In this test-suite directory, the central region of the model where the entropy per baryon reaches 4 , is first removed. This excision represents the part of the model that would have collapsed to form a proto-neutron star. The model is further allowed to continue infalling until its inner boundary (IB) reaches close to the location of the stalled shock. After the first few seconds, we account for further fallback by removing negative velocity material at the IB. The explosion is simulated by injecting energy in a thin layer of approximately 0.2 M⊙ at the IB for 0.2 seconds, at a rate sufficient enough to raise the total energy of the model to our specified value of . We also set the amount of nickel () synthesized in the SN. The mass of the SN ejecta () is determined by subtracting the final core mass after fallback () from the mass of the progenitor model at pre-SN stage (). Thus, = - . The detailed explanation of how the models are simulated from Fe-core infall till near SBO stages is documented in Section 6.1 of B. Paxton et al. (2018).
@@ 3.1 Constraints on , , and
Once the models have reached the near SBO stages with desired and , we utilize the radiation hydrodynamics code, STELLA (S. I. Blinnikov et al., 1998; S. Blinnikov et al., 2000; S. Blinnikov & E. Sorokina, 2004; S. I. Blinnikov et al., 2006) to compute the bolometric light curve and photospheric velocity evolution. STELLA numerically solves the radiative transfer equations by employing the intensity moment approximation in each frequency bin. In typical MESA applications, it operates with 40 frequency groups, sufficient to generate the spectral energy distributions.
The left-hand side of the top panel of Figure 3 shows the comparison of the STELLA-generated bolometric luminosity light curves from the synthetic explosions of our models with the observed bolometric luminosity light curve, while the top right-hand panel shows the comparison of the STELLA-generated photospheric velocity () evolution with the observed Fe II (5169 Å)-line velocities of SN 2024ggi. The estimated by STELLA at a particular instant is the velocity at the location of the photosphere where the Rosseland optical depth () is 2/3 (B. Paxton et al., 2018). The observed bolometric light curve and Fe II-line velocities are adopted from K. Ertini et al. (2025). The most plausible progenitor having an initial mass of 11 M⊙ requires = 0.71051 erg and = 0.055 M⊙ to iterate the observed bolometric light curve (barring the initial 30 days, details about it in the next section) and Fe II velocities. A recent study by L. Dessart et al. (2025) indicates a slightly higher of 0.06 M⊙. However, it is consistent with the amount (0.05–0.06 M⊙) obtained by L. Ferrari et al. (2025). The in our modeling seems to be lower than W. V. Jacobson-Galán et al. (2024); K. Ertini et al. (2025); however, it is in line with the delayed neutrino explosion mechanism of CCSNe as investigated in N. Soker (2024).
@@
Now, more than a year since the discovery of SN 2024ggi, a few nebular phase spectroscopic observations are also there in the literature. Comparing the molecule-free radiative-transfer model-generated spectra with nebular phase spectroscopy, L. Dessart et al. (2025) constrain an explosion energy and nickel mass of 1051 erg and 0.06 M⊙, respectively. Further, while studying the late-time bolometric light curve, L. Ferrari et al. (2025) propose a nickel mass of 0.05 – 0.06 M⊙ for SN 2024ggi. The comparisons of the [O I]/[Ca II] flux ratio in the nebular phase spectra with models from A. Jerkstrand et al. (2014) and L. Dessart et al. (2021) reveal a progenitor having initial mass of around 10–12 M⊙ (L. Ferrari et al., 2025). These observational constraints on nickel mass and initial mass are well in line with our analysis.
3.2 Constraints on the CSM properties
@@
We see in the previous section that the bolometric luminosities of the initial few days ( 30 days) can not be reproduced by any of the models. An ad-hoc mechanism of CSM interaction has been proposed to account for the additional luminosities during the initial days of such Type II SNe. The presence of CSM around the SN 2024ggi progenitor has been confirmed by several observations (W. V. Jacobson-Galán et al., 2024; T. Pessi et al., 2024; M. Shrestha et al., 2024; J. Zhang et al., 2024; T.-W. Chen et al., 2025).
@@
To model the CSM around the possible progenitor of SN 2024ggi, we assume steady state wind with velocity, = 37 km s-1. The choice of this velocity is supported by the value of 374 km s-1 obtained by M. Shrestha et al. (2024) using high-resolution spectroscopic observations during the early phases of SN 2024ggi. For all the models with CSM, we incorporate the wind duration, = 0.5 yr and mass-loss rate, = 1.0 M⊙ yr-1. With the assumed mass-loss rate, the amount of CSM mass () around our progenitor models is 0.5 M⊙, which is slightly higher than T.-W. Chen et al. (2025) but similar to K. Ertini et al. (2025). For our preferred model having initial mass of 11 M⊙, the extent of CSM reaches up to 1.21014 cm ( 1660 R⊙). The inferred mass-loss indicates that the progenitor suffered an enhanced mass-loss during the last 180 days before the explosion. We find that the in our modeling to produce the right amount of CSM responsible for ejecta-CSM interaction is smaller than = 3.6 M⊙ yr-1 obtained by K. Ertini et al. (2025) in the steady wind case. The explosion results with steady-wind CSM are shown in the middle panels of Figure 3. It may not be very likely for any RSG star to have such a high mass-loss rate. As a consequence, we also incorporate the accelerated-wind (also know-wind) scenario as prescribed by T. J. Moriya et al. (2018).
@@
For the -wind case, we adopt the velocity law for the wind as mentioned in T. J. Moriya et al. (2018), . Here, is the initial wind velocity, and it is so chosen that the CSM density is smoothly connected from the surface of the progenitor. is the radial coordinate where the CSM is attached to the progenitor model. is the terminal wind velocity which is 37 km s-1 from M. Shrestha et al. (2024). We also use =0.1 km s-1 (from K. Ertini et al. 2025) and = 4.5 (accepted values being 1–5, as presented in T. J. Moriya et al. 2018) while modeling the initial stages of SN 2024ggi. We tried different combination of CSM properties, however, the preferred model having initial mass of 11 M⊙ with accelerated-wind scenario required a mass-loss rate = 110-2 M⊙ yr-1 with the CSM extended upto 1.11014 cm ( 1600 R⊙). The bottom panels of Figure 3 show the corresponding comparisons of the bolometric light curve (left) and photospheric velocity evolution (right) for the -wind CSM case. The accelerated-wind mass-loss rate in our case is in line with W. V. Jacobson-Galán et al. (2024); M. Shrestha et al. (2024); J. Zhang et al. (2024). An amount of 0.7 M⊙ is required to explain the initial enhancement in the luminosity. Thus, the accelerated-wind CSM structure requires a slightly higher amount and nearly similar extent of CSM to explain the early enhancement in the luminosity. From the velocity evolution panels in the second column of the Figure 3, we notice that at the early times, the favored models (i.e., the models with 11 M⊙) show little/no difference between the photospheric velocity and that of the Fe II line. However, as the photosphere moves deeper into the ejecta, we see that the two velocities start to diverge substantially and the actual measured Fe II velocity point lies slightly higher than the curve (a well-known behavior mentioned in B. Paxton et al., 2018).
@@
The effect of changing on the light curve is shown in the bottom panels of Figure 3. A small increment in (0.1 to 0.2) has a negligible effect on the light curve shape; however, a large change (0.1 to 1.0) significantly reduces the plateau duration. Due to an increased mass-loss rate, this reduction in plateau duration is attributed to the reduced amount of hydrogen in the outer envelope, which controls plateau duration. This analysis also diminishes the possibility of even lower (say, 9 M⊙) initial mass progenitor for SN 2024ggi, as it might not have massive enough hydrogen envelope to explain the observed plateau duration.
@@
To provide a detailed view of the CSM assumed in the present analysis, the density structures of the steady-wind and accelerated-wind scenarios are shown in the left-hand panel of Figure 4. The right-hand panel of Figure 4 displays the wind velocity distribution for the two CSM structures. In case of -wind CSM, the wind velocity near the progenitor star is very small, which smoothly increases as the radius increases; however, in the steady state CSM case, the wind velocity remains fixed to 37 km s-1 throughout.
@@
From the preferred 11 M⊙ model with accelerated-wind CSM, we try to constrain the and by varying one of these two parameters while keeping the other fixed. As shown in Figure 5, a reduced of 0.61051 erg seems to over produce the plateau duration in the observed bolometric luminosity light curve. If we increase the to 0.81051 erg, the tail phase further improves, although one can notice that the plateau luminosity is increasing. Thus, we can not vary the explosion energy much and provide a range of [0.7–0.8]1051 erg. Further, we see that the model produced tail phase luminosity with = 0.065 M⊙ is certainly over-estimated. Thus, due to the absence of a robust tail-phase follow-up observations, an upper limit of 0.055 M⊙ is provided for . This limit on aligns perfectly with the values quoted with recent nebular phase spectroscopy and tail phase analysis of SN 2024ggi (L. Ferrari et al., 2025; L. Dessart et al., 2025).
@@ 4 Discussion and Conclusion
Among all the models in our study, the location of 11 M⊙ at its pre-SN stage on the HR-diagram coincided well with the directly detected progenitor in pre-explosion images. However, we found it difficult to completely rule out the 10 M⊙ and 12 M⊙ models due to their proximity to the directly detected progenitor on the HR-diagram. Thus, although the 11 M⊙ progenitor appeared more plausible, we provided a constrain of 11 M⊙ on the progenitor mass. The pre-SN radius of the preferred 11 M⊙ model was 800 R⊙, however, due to the mass constraint obtained, the range of pre-SN radius would be [690 – 900] R⊙. These constraints on mass and radius were in line with the observationally estimated values reported in D. Xiang et al. (2024); X. Hong et al. (2024).
To explain the overall observed bolometric light curve and velocity evolution, the plausible model with an initial mass of 11 M⊙ required an of [0.7–0.8]1051 erg, of 0.055 M⊙, of 9.1 M⊙, and an amount of 0.7 M⊙ of CSM extended up to 1600 R⊙ (assuming accelerated-wind CSM). However, due to the constraint of 11 M⊙ on the initial mass, the range of pre-SN radius and ejecta mass would be [690–900] R⊙, and [8.2–9.6] M⊙. The overall pre-SN properties and explosion results for 13 and 14 M⊙ models in our simulations disfavored a high mass progenitor for SN 2024ggi.
Comment on the initial mass degeneracy: While performing hydrodynamic modeling of the Type IIP SNe, J. A. Goldberg & L. Bildsten (2020) indicated initial mass degeneracies. Different initial mass progenitors could produce similar bolometric light curves and velocity evolution from different initial physical conditions (e.g., rotation, mass-loss, etc.). However, we saw how changing the initial parameters (rotation and eta) in our models impacted their overall evolution and terminating positions on the HR-diagram. In the present study, we had additional informations on the progenitor’s luminosity, effective temperature, and mass through pre-explosion detection. Thus, the preferred initial mass would be the one that lied in the vicinity of the position of the pre-explosion detected progenitor on the HR diagram. Ultimately, our constrained progenitor with initial mass of 11 M⊙ not only matched the position of the pre-explosion detected progenitor on the HR-diagram, but also explained the observed bolometric light curve and velocity evolution satisfactorily.

### v2

@@ 2.3 Constrain on the physical properties of the possible progenitor
The evolutionary tracks of the models shown in Figure 2 start from ZAMS and terminate at their respective pre-SN stages marked by for each model. The position of the possible progenitor detected in the pre-explosion images from HST (results from 2024ApJ...969L..15X) is also indicated. The effective temperature , and luminosity log(/L⊙) of the 11 M⊙ model at pre-SN stage seems to align well with the progenitor candidate directly detected from HST in archival images. At the pre-SN stage, the 11 M⊙ model has a radius of 800 R⊙. Additionally, the age (log[/yr]) of the 11 M⊙ model at the pre-SN stage is 7.38, matching closely with 2024ApJ...977L..50H.
Although the pre-SN properties of 11 M⊙ model seems to match the physical properties of the possible progenitor detected in pre-explosion images, it is difficult to completely rule out the 10 M⊙ and 12 M⊙ models. These models have pre-SN radii of 690 R⊙ and 900 R⊙, respectively. Thus, for the progenitor of SN 2024ggi, we provide a constrain of 11 M⊙ on the initial mass, and a range of pre-SN radius of 690 R⊙–900 R⊙. These constraints align nicely with the values obtained by analyzing the pre-explosion images (as in 2024ApJ...969L..15X; 2024ApJ...977L..50H). The error bar on our initial mass constraint represents the range of masses lying in the closest vicinity of the error bar of the pre-explosion detected progenitor on the HR-diagram. The details of the physical properties of the models at different stages are provided in Table 1.
@@ 3 Constraints on the explosion properties
In the previous section, we explored the physical properties of the models at their pre-SN stages marked by the onset of Fe-core infall. Next, the models are allowed to continue evolving until they reach near-shock breakout (SBO) stages utilizing the ccsn_IIp test-suite directory of MESA. The pre-SN model from the 12M_pre_ms_to_core_collapse directory is incorporated as the starting point in the ccsn_IIp test-suite directory. In this test-suite directory, the central region of the model where the entropy per baryon reaches 4 , is first removed. This excision represents the part of the model that would have collapsed to form a proto-neutron star. The model is further allowed to continue infalling until its inner boundary (IB) reaches close to the location of the stalled shock. After the first few seconds, we account for further fallback by removing negative velocity material at the IB. The explosion is simulated by injecting energy in a thin layer of approximately 0.2 M⊙ at the IB for 0.2 seconds, at a rate sufficient enough to raise the total energy of the model to our specified value of . We also set the amount of nickel () synthesized in the SN. The mass of the SN ejecta () is determined by subtracting the final core mass after fallback () from the mass of the progenitor model at pre-SN stage (). Thus, = - . The detailed explanation of how the models are simulated from Fe-core infall till near SBO stages is documented in Section 6.1 of 2018ApJS..234...34P.
@@ 3.1 Constraints on , , and
Once the models have reached the near SBO stages with desired and , we utilize the radiation hydrodynamics code, STELLA (1998ApJ...496..454B; 2000ApJ...532.1132B; 2004Ap&SS.290...13B; 2006A&A...453..229B) to compute the bolometric light curve and photospheric velocity evolution. STELLA numerically solves the radiative transfer equations by employing the intensity moment approximation in each frequency bin. In typical MESA applications, it operates with 40 frequency groups, sufficient to generate the spectral energy distributions.
The left-hand side of the top panel of Figure 3 shows the comparison of the STELLA-generated bolometric luminosity light curves from the synthetic explosions of our models with the observed bolometric luminosity light curve, while the top right-hand panel shows the comparison of the STELLA-generated photospheric velocity () evolution with the observed Fe II (5169 Å)-line velocities of SN 2024ggi. The estimated by STELLA at a particular instant is the velocity at the location of the photosphere where the Rosseland optical depth () is 2/3 (2018ApJS..234...34P). The observed bolometric light curve and Fe II-line velocities are adopted from 2025arXiv250301577E. The most plausible progenitor having an initial mass of 11 M⊙ requires = 0.71051 erg and = 0.049 M⊙ to iterate the observed bolometric light curve (barring the initial 30 days, details about it in the next section) and Fe II velocities. We provide a discussion on the choice of this nickel mass and the consequence of choosing a lower or higher nickel mass in the next section. A recent study by 2025arXiv250705803D indicates a slightly larger of 0.06 M⊙. However, it is close to the amount (0.05–0.06 M⊙) obtained by 2025arXiv250722794F. The in our modeling seems to be smaller than 2024ApJ...972..177J; 2025arXiv250301577E; however, it is in line with the delayed neutrino explosion mechanism of CCSNe as investigated in 2024OJAp....7E..69S.
@@
Now, more than a year since the discovery of SN 2024ggi, a few nebular phase spectroscopic observations are also there in the literature. Comparing the molecule-free radiative-transfer model-generated spectra with nebular phase spectroscopy, 2025arXiv250705803D constrain an explosion energy and nickel mass of 1051 erg and 0.06 M⊙, respectively. Further, while studying the late-time bolometric light curve, 2025arXiv250722794F propose a nickel mass of 0.05 – 0.06 M⊙ for SN 2024ggi. The comparisons of the [O I]/[Ca II] flux ratio in the nebular phase spectra with models from 2014MNRAS.439.3694J and 2021A&A...652A..64D reveal a progenitor having initial mass of around 10–12 M⊙ (2025arXiv250722794F). These observational constraints on nickel mass and initial mass are well in line with our results derived from simulations.
3.2 Constraints on the CSM properties and the effect of variation of and
@@
We see in the previous section that the bolometric luminosities of the initial few days ( 30 days) can not be reproduced by any of the models. An ad-hoc mechanism of CSM interaction has been proposed to account for the additional luminosities during the initial days of such Type II SNe. The presence of CSM around the SN 2024ggi progenitor has been confirmed by several observations (2024ApJ...972..177J; 2024A&A...688L..28P; 2024ApJ...972L..15S; 2024ApJ...970L..18Z; 2025ApJ...983...86C).
@@
To model the CSM around the possible progenitor of SN 2024ggi, we assume steady state wind with velocity, = 37 km s-1. The choice of this velocity is supported by the value of 374 km s-1 obtained by 2024ApJ...972L..15S using high-resolution spectroscopic observations during the early phases of SN 2024ggi. For all the models with CSM, we incorporate the wind duration, = 0.5 yr and mass-loss rate, = 1.0 M⊙ yr-1. With the assumed mass-loss rate, the amount of CSM mass () around our progenitor models is 0.5 M⊙, which is slightly higher than 2025ApJ...983...86C but similar to 2025arXiv250301577E. For our preferred model having initial mass of 11 M⊙, the extent of CSM reaches up to 1.21014 cm ( 1660 R⊙). The inferred mass-loss indicates that the progenitor suffered an enhanced mass-loss during the last 180 days before the explosion. We find that the in our modeling to produce the right amount of CSM responsible for ejecta-CSM interaction is smaller than = 3.6 M⊙ yr-1 obtained by 2025arXiv250301577E in the steady wind case. The explosion results with steady-wind CSM are shown in the middle panels of Figure 3. It may not be very likely for any RSG star to have such a high mass-loss rate. As a consequence, we also incorporate the accelerated-wind (also known as the -wind) scenario as prescribed by 2018MNRAS.476.2840M.
@@
For the accelerated-wind case, we adopt the velocity law for the wind as mentioned in 2018MNRAS.476.2840M, . Here, is the initial wind velocity, and it is so chosen that the CSM density is smoothly connected from the surface of the progenitor. is the radial coordinate where the CSM is attached to the progenitor model. is the terminal wind velocity which is 37 km s-1 from 2024ApJ...972L..15S. We also use = 0.1 km s-1 (from 2025arXiv250301577E) and = 4.5 (accepted values being 1–5, as presented in 2018MNRAS.476.2840M) while modeling the initial stages of SN 2024ggi. We tried different combination of CSM properties, however, the preferred model having initial mass of 11 M⊙ with accelerated-wind scenario required a mass-loss rate = 110-2 M⊙ yr-1 with the CSM extended upto 1.11014 cm ( 1600 R⊙). The bottom panels of Figure 3 show the corresponding comparisons of the bolometric light curve (left) and photospheric velocity evolution (right) for the -wind CSM case. The accelerated-wind mass-loss rate in our case is in line with 2024ApJ...972..177J; 2024ApJ...972L..15S; 2024ApJ...970L..18Z. An amount of 0.7 M⊙ is required to explain the initial enhancement in the luminosity. Thus, the accelerated-wind CSM structure requires a slightly higher amount and nearly similar extent of CSM to explain the early enhancement in the luminosity. From the velocity evolution panels in the second column of the Figure 3, we notice that at the early times, the favored models (i.e., the models with 11 M⊙) show little/no difference between the photospheric velocity and the Fe II line velocities. However, as the photosphere moves deeper into the ejecta, we see that the two velocities start to diverge substantially and the actual measured Fe II velocity point lies slightly higher than the curve (a well-known behavior mentioned in 2018ApJS..234...34P).
@@
The effect of changing on the light curve is also shown in the bottom panels of Figure 3. A small increment in (0.1 to 0.2) has a negligible effect on the light curve shape; however, a large change (0.1 to 1.0) significantly reduces the plateau duration. Due to an increased mass-loss rate, this reduction in plateau duration is attributed to a smaller outer hydrogen envelope, which controls plateau duration. This analysis also rules out the possibility of a much lower (say, 9 M⊙) initial mass progenitor for SN 2024ggi, as it might not have massive enough hydrogen envelope to explain the observed plateau duration.
@@
To provide a detailed view of the CSM assumed in the present analysis, the density structures of the steady-wind and accelerated-wind scenarios are shown in the left-hand panel of Figure 4. The right-hand panel of Figure 4 displays the wind velocity distribution for the two CSM structures. In case of -wind CSM, the wind velocity near the progenitor star is very small, which smoothly increases as the radius increases; however, in the steady state CSM case, the wind velocity remains fixed to 37 km s-1 throughout. A recent study by 2025Galax..13...72V pointed out that RSGs have extended, dense atmospheres from which only a fraction is lost through a stellar wind. The high inferred mass-loss rates in the years preceding an SN explosion likely reflect the dense atmosphere, not a wind, and therefore do not represent mass loss and should instead by quantified in terms of density only.
@@
From the preferred 11 M⊙ model with accelerated-wind CSM, we try to constrain the and by varying one of these two parameters while keeping the other fixed. As shown in Figure 5, a reduced of 0.61051 erg seems to slightly over produce the plateau duration in the observed bolometric luminosity light curve. If we increase the to 0.81051 erg, the tail phase improves, although one can notice that the plateau luminosity is increasing. Thus, we can not vary the explosion energy much and provide a range of [0.7–0.8]1051 erg. As shown in Fig. 5, the model with of 0.049 M⊙ nicely iterates the observed luminosities. The amount of the nickel mass absolutely matches to the value estimated by the analytical model fit represented by the equation 1 in 2025arXiv250722794F. Further, we see that the models with = 0.055 M⊙ and 0.065 M⊙ slightly over-estimate the nebular phase luminosity. Moreover, a lesser amount (0.035 M⊙ and 0.045 M⊙) of underestimates the nebular phase luminosity. Thus, we robustly constrain an amount of 0.049 M⊙ of nickel synthesized in SN 2024ggi.
@@ 4 Discussion and Conclusion
Among all the models in our study, the location of 11 M⊙ at its pre-SN stage on the HR-diagram matched well with the directly detected progenitor candidate in the pre-explosion images. However, we found it difficult to completely rule out the 10 M⊙ and 12 M⊙ models due to their proximity to the directly detected progenitor candidate on the HR-diagram. Thus, although the 11 M⊙ progenitor appeared more plausible, we provided a constrain of 11 M⊙ on the progenitor mass. The pre-SN radius of the preferred 11 M⊙ model was 800 R⊙, however, due to the mass constraint obtained, the range of pre-SN radius would be [690 – 900] R⊙. These constraints on mass and radius were in line with the observationally estimated values reported in 2024ApJ...969L..15X; 2024ApJ...977L..50H.
To explain the overall observed bolometric light curve and velocity evolution, the plausible model with an initial mass of 11 M⊙ required an of [0.7–0.8]1051 erg, of 0.049 M⊙, of 9.1 M⊙, and an amount of 0.7 M⊙ of CSM extended up to 1600 R⊙ (assuming accelerated-wind CSM). However, due to the constraint of 11 M⊙ on the initial mass, the range of pre-SN radius and ejecta mass would be [690–900] R⊙, and [8.2–9.6] M⊙. The overall pre-SN properties and explosion results for 13 and 14 M⊙ models in our simulations disfavored a high mass progenitor for SN 2024ggi.
Comment on the initial mass degeneracy: While performing hydrodynamic modeling of the Type IIP SNe, 2020ApJ...895L..45G indicated initial mass degeneracies. Different initial mass progenitors could produce similar bolometric light curves and velocity evolution from different initial physical conditions (e.g., rotation, mass-loss, etc.). However, we saw how changing the initial parameters (rotation and eta) in our models impacted their overall evolution and terminating positions on the HR-diagram. In the present study, we had additional informations on the progenitor candidate’s luminosity, effective temperature, and mass through pre-explosion progenitor candidate detection. Thus, the preferred initial mass would be the one that lied in the proximity of the position of the pre-explosion detected progenitor candidate on the HR diagram. Ultimately, our constrained progenitor with initial mass of 11 M⊙ not only matched the position of the pre-explosion detected progenitor candidate on the HR-diagram, but also explained the observed bolometric light curve and velocity evolution satisfactorily.